OpenStudy (anonymous):

Solve 2tan^2 x + tanx - 3 = 0

OpenStudy (anonymous):

Ok lets start off with subsituting x=tanx

OpenStudy (anonymous):

\[2x^2+x-3=0\] Now we have a quadratic equation and we must solve for x

OpenStudy (anonymous):

(2tan+3)(tan-1)=0

OpenStudy (anonymous):

\[2x^2-2x+3x-3=0\] \[2x(x-1)+3(x-1)=0\] \[(2x+3)(x-1)=0\] \[ x= - \frac{3}{2} \text{ and }x=1\]

OpenStudy (anonymous):

Now let us plug tan x back into our equation \[\tan x= - \frac{3}{2} \text { and } \tan x=1\] \[ x= \tan^{-1}(- \frac{3}{2}) \text { and } x=\tan^{-1}(1)\]

OpenStudy (anonymous):

@christen_j0 Im leaving the rest for you to solve